Solve the problem and show the complete solution.

A marine weather station detects waves 9.28 m long and 1.65 m high and travels a distance of
50.0 m in 21.8 seconds. Determine the speed and the frequency of the waves.
1
Expert's answer
2020-12-17T09:13:23-0500
"v=\frac{s}{t}=\frac{50}{21.8}=2.29\frac{m}{s}\\f_1=\frac{v}{\lambda_1}=\frac{2.29}{9.28}=0.247\ Hz\\f_2=\frac{v}{\lambda_2}=\frac{2.29}{1.65}=1.39\ Hz"